OpenStudy (anonymous):

Tiara wants to divide a square piece of paper into two equalivalent triangles. If the square measures 20 cm on each side, what will the third side of each triangle measure?

OpenStudy (anonymous):

\[20\sqrt{2}\]

OpenStudy (pfenn1):

Use the equation a^2 + b^2 = c^2

OpenStudy (anonymous):

length of the diagonal of the square

OpenStudy (pfenn1):

@sheg, we should be trying to help them figure it out, not just give the answer

OpenStudy (anonymous):

@pfenn1 yes buddy as you wish i would try to resolve the issue rather than shooting the answer

OpenStudy (pfenn1):

Sorry, @sheg. I don't mean to slap your wrist. But I really want to help learning, not just finishing the question. I am sure that folks are very grateful to you because you are very good at getting to the core of the problem. Much more succinct than I am.

OpenStudy (anonymous):

@pfenn1 buddy never mind actually i too start giving hints towards the answers but someone comes in between and throws the answer so i too threw the answer before anyone else

OpenStudy (pfenn1):

@sheg, yes, that does happen. I'm still pretty new here so am still trying to get an idea of how to proceed.

OpenStudy (anonymous):

i am not new actually i logged in back after a long time @pfenn1

OpenStudy (pfenn1):

see you around @sheg